A cashier has 30 bills, all of which are $10 or $20 bills. The total value of the money is $330. How many of each type does the cashier have?
Answer by J2R2R(94)

A cashier has 30 bills, all of which are $10 or $20 bills. The total value of the money is $330. How many of each type does the cashier have?

Let the number of $10 bills be m and the number of $20 bills be n.

We know m + n = 30
and 10m + 20n = 330 (m $10 bills + n $20 bills totalling $330).

So we have two simultaneous equations with the third equation a multiple (10) of the first:
(1) m + n = 30
(2) 10m + 20n = 330
(3) 10m + 10n = 300

The difference of last two equations
10n = 30
Therefore n = 3 and m = 27
27 x $10 + 3 * $20 = $330 using 30 bills.